1.A firm is considering two mutually exclusive projects, X and Y, with the following cash flows:

0 1 2 3 4
Project X -$1,000 $90 $320 $430 $700
Project Y -$1,000 $1,000 $90 $45 $45

The projects are equally risky, and their WACC is 13%. What is the MIRR of the project that maximizes shareholder value? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

X:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=90/1.13+320/1.13^2+430/1.13^3+700/1.13^4

=$1057.59

NPV=Present value of inflows-Present value of outflows

=1057.59-1000

=$57.59(Approx).

Y:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=1000/1.13+90/1.13^2+45/1.13^3+45/1.13^4

=$1014.23

NPV=Present value of inflows-Present value of outflows

=1014.23-1000

=$14.23(Approx).

Hence X is better having higher NPV.

Hence:

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value of inflows=90*(1.13)^3+320*(1.13)^2+430*(1.13)+700

=$1724.36873

MIRR=[Future value of inflows/Present value of outflows]^(1/time period)-1

=[1724.36873/1000]^(1/4)-1

=14.59%(Approx).